OpenStudy (anonymous):

find an equation of the line that is tangent to the graph of the function f(x)=7/(√x) and parallel to the line 7x+2y-6=0

OpenStudy (raden):

first, can you determine the slope of line 7x+2y-6=0 ?

OpenStudy (anonymous):

@raden 7/2?

OpenStudy (raden):

the slope of line ax + by + c = 0 is m = -a/b (with m is slope) see the equation 7x+2y-6=0, it means a = 7, b = 2 therefore m = -a/b = -7/2

OpenStudy (anonymous):

ok

OpenStudy (raden):

then we knowed that 2 lines are parallel if the slope is same. so use m = -7/2 too

OpenStudy (raden):

now find derivative of f(x) = 7/(√x) = 7/x^1/2 = 7x^(-1/2) f ' (x) = -1/2 * 7 * x^-3/2 = -7/2 x^(-3/2) = -7/(2x^3/2) because the slope (m) = -7/2, then -7/(2x^3/2) = -7/2 solve for x = ........ ?

OpenStudy (anonymous):

I got x=1 but thats not an answer choice. ANSWER CHOICES: 3x+y-6=0 7x+2y-14=0 7x+2y-21=0 3x+2y+3=0 3x+y+21=0

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

yeah, it just point slope coordinat, not tangent equation :) now subtitute x = 1 into f(x)=7/(√x) to get y

OpenStudy (raden):

y = 7/(√1) = 7,\ so it is (1, 7) as the point slope

OpenStudy (anonymous):

ok

OpenStudy (raden):

finally, use the formula of tangent line : y - y1 = m (x - x1) y - 7 = -7/2 (x - 1) times 2 on both sides, we get 2(y - 7) = -7(x - 1) 2y - 14 = -7x + 7 add 7x on both sides : 7x + 2y - 14 = 7x -7x + 7 7x + 2y - 14 = 7 subtracted by 7, on both sides : 7x + 2y - 14 - 7 = 7 - 7 7x + 2y - 21 = 0

OpenStudy (anonymous):

thanks!

OpenStudy (raden):

you're welcome ;)